# Speedups in nonequilibrium thermal relaxation: Mpemba and related effects

### v1

@@ 5.3 Outlook on statistics of the Mpemba effect
For the ensembles considered above, the Mpemba effect occurs with probabilities $\gtrsim 0.35$, suggesting that the Mpemba effect likely does not have “self-averaging properties”: the properties of one large system do not determine the typical realization. Even in large system sizes, different realizations have different Mpemba properties. This fact encourages careful design of relaxation protocols, given an energy landscape that could, in principle, lead to faster relaxation. The REM was introduced as a simple disordered model that shares several qualitative features with spin-glasses. The Mpemba effect has been studied for spin-glasses [59] (see Sec. 7.1), where the connection with the correlation length has been observed. It would be interesting to study the statistics of the Mpemba effect in spin-glass systems and compare it with the results obtained on the REM.
Besides annealing, cooling, and heating protocols, questions of efficient relaxation also arise in a variety of contexts, such as the design of efficient samplers. In particular, Markov chain Monte Carlo algorithms use Markov chains to create realizations of the system at its steady state [161, 162]. It is often desirable to accelerate the relaxation of a Markov chain to a target steady state [163, 164]. One possible way to do so could be to design relaxation dynamics (barriers) that would show the Mpemba effect for a given system and initial condition. For example, although there is no Mpemba effect for a complete graph with Metropolis dynamics, adding barriers that remove some state transitions can induce a Mpemba effect [96]. We discuss more Markov chain Monte Carlo sampling in connection to the Mpemba effect in Sec. 8.2.5.
@@
We should also mention the study of Meibohm and Knapp on the asymptotic relaxation rate of a system perturbed by quenched disorder [166]. Averaging over disorder, they found that the mean and variance of the asymptotic relaxation rate $\lambda_{2}$ depends nonmonotonically on the parameters of a broad class of disorder realizations. Thus, the type of disorder affects the asymptotic relaxation rate. It would be interesting to consider further the statistics of spectral changes to the relaxation rate matrix (operator) given such perturbations. And especially to explore them in the context of potential relaxation speedup with quenched disorder.
@@ 6 Experiments
In the Introduction (Section 1), we reviewed historical observations of anomalous behavior in the freezing of water, where, after a sudden temperature quench, hot water is claimed to cool or start freezing more quickly than water at intermediate temperatures. Modern discussions of this phenomenon date from the experiments by Mpemba and Osborne in 1969 [1] and the nearly simultaneous, independent work of Kell [26]. We reviewed the complicated history of studies on cooling and freezing water, which continue to this day. However, although water has played an important historical role in suggesting and providing some evidence for the Mpemba effect, the analysis of individual experiments is complicated. More importantly, focusing on one system raises the question of the generality of anomalous cooling (and heating) phenomena. If the time to reach thermal equilibrium is not a monotonic function of the temperature difference between initial and final states, is such behavior observed more generally or only in water? Thus, in this section, we review experiments in other systems.
We begin in Section 6.1 with a survey of observations of Mpemba effects in condensed matter systems other than water. These are important because they were the first to suggest that an explanation of the Mpemba effect should not be based too firmly on the peculiar features of water. Unfortunately, the physics involved is at least as complex as in the water-based experiments and is often more poorly understood. In Section 6.2, we present a set of recent experiments using a colloidal particle diffusing in water and moving in an external potential. The experiments were the first to clearly demonstrate the weak (original) Mpemba scenario, albeit in a simple system. They also provided the first evidence for more subtle phenomena, such as the strong and inverse Mpemba effects (Sections 2.1.6 and 2.1.5). Section 6.3 presents the first experiments to demonstrate the Mpemba effect in quantum systems.
@@ 6.1 Other condensed-matter systems
  • Clathrates. Ahn et al. observed the Mpemba effect experimentally in clathrate hydrates consisting of hydrogen-bonded water frameworks and enclathrated tetrahydrofuran (THF) molecules [134]. They observed that THF solutions display Mpemba-like behavior in the temperature range 278–318 K, where warmer solutions form hydrates faster than colder solutions. At an initial temperature above 318 K, the formation time was delayed, as some THF molecules in gas form in solution at high temperature must condense before they can form hydrates. Ahn et al. defined the freezing temperature as the temperature at which the THF solution structurally transforms into THF hydrates and measured it via differential scanning calorimetry (DSC). They checked that they could observe a Mpemba effect in a pure-water system using DSC and found a nonmonotonic (roughly parabolic) dependence of time to freeze on the initial temperature $T_{i}$ that is characteristic of the Mpemba effect. THF solutions also showed a nonmonotonic relationship between time to freeze and $T_{i}$ (roughly a cubic). Ahn et al. then studied the cooperative relationship between the intramolecular polar-covalent bonds (O–H) and intermolecular hydrogen bonds (O:H) for the THF system. Using Raman spectra, they observed that O:H stretching phonons of water in THF showed a blue shift (stiffened) when cooled, whereas the O–H stretching mode showed a redshift (softened). Such trends are also found in pure water and thus at least qualitatively support a role for hydrogen bonding in anomalous cooling, as suggested by Zhang et al. [45, 46].
@@
  • Magnetic system. Chaddah et al. studied the magnetic—not thermal—relaxation of polycrystalline manganite La0.5Ca0.5MnO3 (LCMO) [171]. A sample was cooled from 320 K to 5 K in different external fields. Then, at 5 K, the field is quenched to a smaller value, and the authors measure the decay of the initial magnetization. They find that states that are initially farther from equilibrium can overtake states that start closer to equilibrium, in a magnetic analog of the Mpemba effect.
@@ 6.2 Colloidal systems
As we shall see, controlling the shape and spatial scale of the potential are important requirements. In the work reported here, the potentials were “virtual potentials” that were created by a feedback loop. The basic procedure is illustrated in Fig. 31 and adapts the feedback (or “ABEL”) trap created for biophysical applications [173] to the study of larger-scale colloidal particles [174]. The Mpemba experiments used a second-generation version of feedback trap that had special hardware to reduce the feedback loop time and applied forces based on the displacements of an optical tweezer (dashed-blue potential in Fig. 31b). The latter was more stable than earlier feedback traps that depended on electrokinetic forces (with attendant chemical drifts).
@@
Using the feedback trap, Kumar and Bechhoefer reported experiments on the standard Mpemba effect and the strong Mpemba effect in Ref. [6]. Then Ref. [64] reported first observations of the inverse Mpemba effect. The general scheme of all of these experiments is to draw an initial condition for a particle from a Boltzmann distribution $\pi(x;T)$ that corresponds to a system in equilibrium at a temperature $T$ that is, in general, different from the bath temperature $T_{b}$. Because the particle moves in water at temperature $T_{b}$, the effective quench from $T$ to $T_{b}$ is instantaneous. This trick of using the initial position of the particle as a substitute for a physical temperature eliminates any complication arising from finite-time transients.
@@
The trajectories $x(t)$ for repeated trials are grouped to form a statistical ensemble, from which, at each time $t$, one can estimate the probability density $p(x,t)$ of beads. This probability density then relaxes to the equilibrium Boltzmann distribution, $\pi(x,T_{b})$. Fig. 32 shows an example of the evolution of the equilibration time as a function of the initial temperature. The time is by definition zero when $T_{\textrm{initial}}=T_{b}$. As $T_{\textrm{initial}}$ increases, the time to equilibrate also increases. Then, for $T_{\textrm{initial}}/T_{b}\approx 10$, there is a local maximum in equilibration time and further increases in initial temperature decrease the equilibration time. This decrease is a signature of the Mpemba effect. The local minimum in equilibration time at $T_{\textrm{initial}}/T_{b}\approx 100$ (indicated by the black arrow) then reflects the strong Mpemba effect. To identify the local minimum with the strong Mpemba effect, the authors used measurements of the relaxation of $p(x,t)$ to equilibrium, to extract the (modulus of the) $a_{2}$ coefficient discussed in Section 3. The strong Mpemba effect corresponds to the vanishing of $|a_{2}|$ and, as a result, an exponential speedup of relaxation.
Figure 32: Equilibration time of ensemble of colloidal particles as a function of the initial temperature, for an asymmetry parameter $\alpha=9$. Arrow indicates the strong Mpemba effect. Source: Figure reproduced with permission from [6].
@@ 6.3 Experiments on quantum Mpemba effects
In January, 2024, three preprints posted to arXiv gave evidence for the Mpemba effect in systems where quantum effects are important. The experimental apparatus in all three works is based on trapped ions, but the three works are quite different: [8] measured an inverse Mpemba effect of the type discussed in Sec.(2.2.1); [9] measured exponential acceleration in the relaxation rate by first applying a unitary transformation on the initial state to annihilate its overlap with the slowest relaxation mode, as discussed in Sec.(2.2.2); and [10] measured the entanglement asymmetry quantum Mpemba effect (discussed in Sec. 2.2.3) in a system of 12 ions.
In all three experiments, the system is composed of one [8, 9] or many [10] ions, trapped in a Paul trap, and laser cooled. This type of setup has become standard, with applications ranging from accurate clocks [175] to precise magnetometry [176] to quantum computing [177]. See [178, 179, 180] for reviews describing the experimental setups. Interestingly, additional experimental manuscript that aims to measure the quantum Mpemba effect was uploaded to the arXiv in January 2024 [181]. However, this manuscript was different from the other three manuscript in two major points: (i) Instead of trapped ions, it used the IBM quantum computer, available through the IBM quantum experience. This computer can be accessed by the general public, and instead of trapped ions it is based on super-conducting Qbits. They considered the single qubit thermal relaxation that exists because of the coupling between the system and its environment, where rather than initializing the qbit in a hot thermal state, it was initialized in a pure state, and the thermal state relaxation trajectory was calculating by a proper averaging over these states. More importantly, (ii) No Mpemba effect, of any type, experimental or theoretical, was observed in this work. We therefore do not discuss this manuscript in what follows.
@@ 6.3.1 Inverse Mpemba effect in a single strontium ion
In [8], a nonequilibrium strong inverse quantum Mpemba effect was demonstrated experimentally on a single trapped 88Sr+ ion. The relevant degrees of freedom in the system are the ion’s internal energy levels, not the motion degrees of freedom. Of all the internal states of the ion, two states are considered as “the qubit system”, whereas other states are used as auxiliary states [8]. The transitions between the two states have several sources: first, there is a coherent driving generated by an oscillating magnetic field with a frequency that is tuned to the specific transition. In addition, there is an incoherent decay transition between the two states, mediated by a thermal bath, which is generated by coupling the system to several auxiliary states in a way that can effectively be analyzed as a thermal bath. Finally, there are dephasing processes, also generated by the thermal bath, namely by the coupling to the auxiliary states. These different processes cannot be coupled simultaneously to the system, so they are done sequentially, i.e., by Trotterization. Because there is coupling both to a thermal bath and to a coherent source, the steady state of the system is not an equilibrium but rather a non-equilibrium steady state. Nonetheless, since coherent driving is kept fixed while the bath temperature varies, the result is an equivalent to the “equilibrium locus” in the classical Mpemba effect.
@@ 6.3.2 Unitary transformation in a single calcium ion
As in the above experiment, in [9] the internal energy states of 40Ca+ ions were used as the relevant degrees of freedom for the relaxation process. Here, three energy levels coupled via lasers having the relevant frequencies. The density matrix describing the system evolves under Lindbladian dynamics as described in Eq. (59), and the steady state of the system is a non-equilibrium steady state, not a Boltzmann distribution.
@@ 6.3.3 Entanglement-asymmetry Mpemba effect in a linear string of twelve calcium ions
Figure 33: Experimental demonstration of the Entanglement asymmetry quantum Mpemba effect. The system (upper inset) is a 1D system of twelve trapped ions, of which four constitute “the system” (subsystem $A$), and the rest the “bath” (subsystem $\tilde{A}$). The Entanglement Asymmetry (EA) of subsystem $A$ changes as a function of a global rotation angle, $\theta$ (lower inset). Initiating the system at different values of $\theta$ and observing the entanglement asymmetry evolving in time shows a clear crossing between the relaxation trajectories. The figure is reproduced from [10].
@@
For a 1D system, it is clear that in the thermodynamic limit any small subsystem should relax to a Boltzmann distribution that would restore the $U(1)$ symmetry after a long time, since there are no phase transitions in the system. Thus, the above entanglement asymmetry should decay as a function of time. Indeed, as shown in Fig. 33, the entanglement asymmetry decays with time for all initial conditions. Moreover, the same figure demonstrates the existence of anomalous relaxation in the system, where initial conditions with higher entanglement asymmetry restore the symmetry faster than initial conditions with lower initial value of the entanglement asymmetry.
@@ 7.1.1 Spin glass system
After the preparation stage, the systems were quenched to the spin glass phase by coupling to a bath at $T_{final}=0.7\approx 0.64\ T_{c}$. As the final temperature is below the glass transition, the relaxation is extremely slow, and even $10^{11}$ simulation sweeps were not enough to observe the final stage of relaxation, as expected in a glassy system. To find a Mpemba effect, the energy per spin (averaged over both trajectories and the quenched disorder $J_{x,y}$) was calculated through the relaxation dynamics. Crossings of the energy relaxation curves versus time were taken to indicate a Mpemba effect. Such crossings, for initial temperatures above and below the glass transition and with subsequent relaxation below the glass transition temperature, can be seen on Figs. 34 and 35. An inverse Mpemba effect was also observed; in this case, all three temperatures were below the glass temperature of the model.
Figure 34: Mpemba effect in the spin-glass system starting from a paramagnetic phase, which is quenched below the critical temperature, $T_{c}=1.102$, to the spin-glass phase, with $T_{final}=0.7\approx 0.64\ T_{c}$. The hot start, ($T_{1}=1.3$, $\xi_{1}=12$, $E_{1}(t=0)=-1.6428$), overtakes the two cold starts ($T_{2}=1.2$, $\xi_{2}=5$, $E_{2}(t=0)=-1.6714$) and ($T_{3}=1.2$, $\xi_{3}=8$, $E_{3}(t=0)=-1.6738$). The correlation length grows for temperatures close to $T_{c}$; thus, the hot start, since it has the longest correlation length, is the “closest” to the state the system is relaxing to. The order of the two crossings implies the importance of the correlation length. The error bars are depicted in the thickness of the lines. The figure is reproduced from [59].
@@
Figure 35: a) Mpemba effect in the spin-glass system below the glass transition, $T_{c}=1.102$, for several correlation lengths. The energy relaxation curves show that the coherence length plays a major role in the relaxation. b) Relationship between energy and coherence length $\xi$. A thermal protocol is responsible for the vertical moves in b) that produce the Mpemba effect. The figures are reproduced from [59].

### v3

@@ 5.3 Outlook on statistics of the Mpemba effect
For the ensembles considered above, the Mpemba effect occurs with probabilities $\gtrsim 0.35$, suggesting that the Mpemba effect likely does not have “self-averaging properties”: the properties of one large system do not determine the typical realization. Even in large system sizes, different realizations have different Mpemba properties. This fact encourages careful design of relaxation protocols, given an energy landscape that could, in principle, lead to faster relaxation. The REM was introduced as a simple disordered model that shares several qualitative features with spin-glasses. The Mpemba effect has been studied for spin-glasses [59] (see Sec. 7.1), where the connection with the correlation length has been observed. It would be interesting to compare the statistics of the Mpemba effect in spin-glass systems and with those found for the REM.
Besides annealing, cooling, and heating protocols, questions of efficient relaxation also arise in other contexts, such as the design of efficient samplers. In particular, Markov chain Monte Carlo algorithms use Markov chains to create realizations of the system at its steady state [170, 171]. It is often desirable to accelerate the relaxation of a Markov chain to a target steady state [172, 173]. One possible approach could be to design relaxation dynamics (barriers) with a Mpemba effect, for a given system and initial condition. For example, although there is no Mpemba effect for a complete graph with Metropolis dynamics, adding barriers that remove some state transitions can induce a Mpemba effect [96]. We discuss more Markov chain Monte Carlo sampling in connection to the Mpemba effect in Sec. 8.2.5.
@@
We should also mention the study of Meibohm and Knapp on the asymptotic relaxation rate of a system perturbed by quenched disorder [175]. Averaging over disorder, they found that the mean and variance of the asymptotic relaxation rate $\lambda_{2}$ depends nonmonotonically on the parameters of a broad class of disorder realizations. Thus, the type of disorder affects the asymptotic relaxation rate. It would be interesting to consider further the statistics of spectral changes to the relaxation rate matrix (operator) given such perturbations. And it would be especially interesting to explore them in the context of a potential relaxation speed up caused by the quenched disorder.
@@ 6 Experiments
In the Introduction (Section 1), we reviewed historical observations of anomalous behavior in the freezing of water, where, after a sudden temperature quench, hot water is claimed to cool or start freezing more quickly than water at intermediate temperatures. Modern discussions of this phenomenon date from the experiments by Mpemba and Osborne in 1969 [1] and the nearly simultaneous, independent work of Kell [26]. We reviewed the complicated history of studies on cooling and freezing water, which continue to this day. However, although water has played an important historical role in suggesting and providing some evidence for the Mpemba effect, the analysis of individual experiments is complicated. More importantly, focusing on one system raises the question of the generality of anomalous cooling (and heating) phenomena. If the time to reach thermal equilibrium is not a monotonic function of the temperature difference between initial and final states, is such behavior observed more generally or only in water? To persuade the reader that the Mpemba effect is not unique to water, we review in this section experiments in other systems.
We begin in Sec. 6.1 with a survey of observations of Mpemba effects in condensed matter systems other than water. These are important because they were the first to suggest that an explanation of the Mpemba effect should not be based too firmly on the peculiar features of water. Unfortunately, the physics involved is at least as complex as in the water-based experiments and is often more poorly understood. In Sec. 6.2, we present a set of recent experiments using a colloidal particle diffusing in water and moving in an external potential. The experiments were the first to clearly demonstrate the weak (original) Mpemba scenario, albeit in a simple system. They also provided the first evidence for more subtle phenomena, such as the strong and inverse Mpemba effects discussed in Sec. 2.1.6 and 2.1.5. Section 6.3 presents the first experiments to demonstrate the Mpemba effect in quantum systems.
@@ 6.1 Other condensed-matter systems
  • 2.
    Clathrates. Ahn et al. observed the Mpemba effect experimentally in clathrate hydrates consisting of hydrogen-bonded water frameworks and enclathrated tetrahydrofuran (THF) molecules [141]. They observed that THF solutions display Mpemba-like behavior in the temperature range 278–318 K, where warmer solutions form hydrates faster than colder solutions. At an initial temperature above 318 K, the formation time was delayed, as some THF molecules in gas form in solution at high temperature must condense before they can form hydrates. Ahn et al. defined the freezing temperature as the temperature at which the THF solution structurally transforms into THF hydrates and measured it via differential scanning calorimetry (DSC). They checked that they could observe a Mpemba effect in a pure-water system using DSC and found a nonmonotonic, roughly parabolic dependence of time to freeze on the initial temperature $T_{i}$ that is characteristic of the Mpemba effect. THF solutions also showed a nonmonotonic relationship between time to freeze and $T_{i}$ (roughly a cubic). Ahn et al. then studied the cooperative relationship between the intramolecular polar-covalent bonds (O–H) and intermolecular hydrogen bonds (O:H) for the THF system. Using Raman spectra, they observed that O:H stretching phonons of water in THF showed a blue shift (stiffened) when cooled, whereas the O–H stretching mode showed a redshift (softened). Such trends are also found in pure water and thus at least qualitatively support a role for hydrogen bonding in anomalous cooling, as suggested by Zhang et al. [45, 46].
  • 3.
@@
  • 5.
    Magnetic system. Chaddah et al. studied the magnetic, not thermal, relaxation of polycrystalline manganite La0.5Ca0.5MnO3 (LCMO) [180]. A sample was cooled from 320 K to 5 K in different external fields. Then, at 5 K, the field is quenched to a smaller value, and the authors measure the decay of the initial magnetization. They find that states that are initially farther from equilibrium can overtake states that start closer to equilibrium, in a magnetic analog of the Mpemba effect.
@@ 6.2 Colloidal systems
As we shall see, controlling the shape and spatial scale of the potential are important requirements. In the work reported here, the potentials were “virtual potentials” that were created by a feedback loop. The basic procedure is illustrated in Fig. 31 and adapts the feedback (or “ABEL”) trap created for biophysical applications [182] to the study of larger-scale colloidal particles [183]. The Mpemba experiments used a second-generation version of feedback trap that had special hardware to reduce the feedback loop time and applied forces based on the displacements of an optical tweezer (dashed-blue potential in Fig. 31b). The latter was more stable than earlier feedback traps that depended on electrokinetic forces and showed attendant chemical drifts.
@@
Using the feedback trap, Kumar and Bechhoefer reported experiments on the standard Mpemba effect and the strong Mpemba effect in Ref. [6]. Then Ref. [64] reported first observations of the inverse Mpemba effect. The procedure for all these experiments is to draw an initial condition for a particle from a Boltzmann distribution $\pi(x;T)$ that corresponds to a system in equilibrium at a temperature $T$ that, in general, differs from the bath temperature $T_{b}$. Because the particle moves in water at temperature $T_{b}$, the effective quench from $T$ to $T_{b}$ is instantaneous. This trick of using the initial position of the particle as a substitute for a physical temperature eliminates any complication arising from finite-time transients.
@@
The trajectories $x(t)$ for repeated trials are grouped to form a statistical ensemble, from which, at each time $t$, one can estimate the probability density $p(x,t)$ of beads. This probability density then relaxes to the equilibrium Boltzmann distribution, $\pi(x,T_{b})$. Figure 32 shows an example of the evolution of the equilibration time as a function of the initial temperature. The time is by definition zero when $T_{\textrm{initial}}=T_{b}$. As $T_{\textrm{initial}}$ increases, the time to equilibrate also increases. Then, for $T_{\textrm{initial}}/T_{b}\approx 10$, there is a local maximum in equilibration time and further increases in initial temperature decrease the equilibration time. This decrease is a signature of the Mpemba effect. The local minimum in equilibration time at $T_{\textrm{initial}}/T_{b}\approx 100$ (indicated by the black arrow) then reflects the strong Mpemba effect. To identify the local minimum with the strong Mpemba effect, the authors used measurements of the relaxation of $p(x,t)$ to equilibrium, to extract the (modulus of the) $a_{2}$ coefficient discussed in Sec. 3. The strong Mpemba effect corresponds to the vanishing of $|a_{2}|$ and, as a result, an exponential speedup of relaxation.
Figure 32: Equilibration time of ensemble of colloidal particles as a function of the initial temperature, for an asymmetry parameter $\alpha=9$. Arrow indicates the strong Mpemba effect. Source: Reprinted with permission from [6].
@@ 6.3 Experiments on quantum Mpemba effects
In January, 2024, three preprints posted to arXiv gave evidence for the Mpemba effect in systems where quantum effects are important. The experimental apparatus in all three works is based on trapped ions, but the three works are quite different: [8] measured an inverse Mpemba effect of the type discussed in Sec. 2.2.1; [9] measured exponential acceleration in the relaxation rate by first applying a unitary transformation on the initial state to annihilate its overlap with the slowest relaxation mode, as discussed in Sec. 2.2.2; and [10] measured the entanglement asymmetry quantum Mpemba effect (discussed in Sec. 2.2.3) in a system of 12 ions.
In all three experiments, the system is composed of one [8, 9] or many [10] ions, trapped in a Paul trap, and laser cooled. This type of setup has become standard, with applications ranging from accurate clocks [184] to precise magnetometry [185] to quantum computing [186]. See [187, 188, 189] for reviews describing the experimental setups. Interestingly, an additional experimental manuscript that aims to measure the quantum Mpemba effect was also uploaded to the arXiv in January 2024 [190]. However, this manuscript differs from the other three manuscripts in two major points: (i) Instead of trapped ions, the authors used a publically available IBM quantum computer that is based on superconducting qubits. The authors considered a single-qubit thermal relaxation that exists because of the coupling between the system and its environment. Rather than initializing the qubit in a hot thermal state, it was initialized in a pure state, and the thermal state relaxation trajectory was calculated by averaging over these states. More importantly, (ii) no Mpemba effect of any type, experimental or theoretical, was observed in this work. We therefore do not discuss this manuscript further.
@@ 6.3.1 Inverse Mpemba effect in a single strontium ion
In [8], a nonequilibrium strong inverse quantum Mpemba effect was demonstrated experimentally on a single trapped 88Sr+ ion. The relevant degrees of freedom in the system are the ion’s internal energy levels, not the motion degrees of freedom. Of all the internal states of the ion, two states are considered as “the qubit system,” whereas other states are used as auxiliary states [8]. The transitions between the two states have several sources: first, there is a coherent driving generated by an oscillating magnetic field with a frequency that is tuned to the specific transition. In addition, there is an incoherent decay transition between the two states, mediated by a thermal bath, which is generated by coupling the system to several auxiliary states in a way that can effectively be analyzed as a thermal bath. Finally, there are dephasing processes, also generated by the thermal bath, namely by the coupling to the auxiliary states. As these different processes cannot be coupled simultaneously to the system, they are done sequentially, i.e., by Trotterization. Because there is coupling both to a thermal bath and to a coherent source, the steady state of the system is not an equilibrium but rather a nonequilibrium steady state. Nonetheless, since coherent driving is kept fixed while the bath temperature varies, the result is an equivalent to the “equilibrium locus” in the classical Mpemba effect.
@@ 6.3.2 Unitary transformation in a single calcium ion
As in the above experiment, in [9] the internal energy states of 40Ca+ ions were used as the relevant degrees of freedom for the relaxation process. Here, three energy levels coupled via lasers having the relevant frequencies. The density matrix describing the system evolves under Lindbladian dynamics as described in Eq. (59), and the steady state of the system is a nonequilibrium steady state, not a Boltzmann distribution.
@@ 6.3.3 Entanglement-asymmetry Mpemba effect in a linear string of twelve calcium ions
Figure 33: Experimental demonstration of the entanglement asymmetry quantum Mpemba effect. The system (upper inset) is a 1D system of twelve trapped ions, of which four constitute “the system” (subsystem $A$), and the rest the “bath” (subsystem $\tilde{A}$). The entanglement asymmetry (EA) of subsystem $A$ changes as a function of a global rotation angle, $\theta$ (lower inset). Initiating the system at different values of $\theta$ and observing the entanglement asymmetry evolving in time shows a clear crossing between the relaxation trajectories. Source: Reprinted with permission from [10].
@@
For a 1D system, it is clear that in the thermodynamic limit any small subsystem should relax to a Boltzmann distribution that would restore the $U(1)$ symmetry after a long time, since there are no phase transitions in the system. Thus, the above entanglement asymmetry should decay as a function of time. Indeed, as shown in Fig. 33, the entanglement asymmetry decays with time for all initial conditions. Moreover, the same figure demonstrates the existence of anomalous relaxation in the system, where initial conditions with higher entanglement asymmetry restore the symmetry faster than initial conditions with lower initial values of the entanglement asymmetry.
@@ 7.1.1 Spin glass system
After the preparation stage, the systems were quenched to the spin-glass phase by coupling to a bath at $T_{final}=0.7\approx 0.64\ T_{c}$. As the final temperature is below the glass transition, the relaxation is extremely slow, and even $10^{11}$ simulation sweeps were not enough to observe the final stage of relaxation, as expected in a glassy system. To find a Mpemba effect, the energy per spin, averaged over both trajectories and the quenched disorder $J_{x,y}$, was calculated through the relaxation dynamics. Crossings of the energy relaxation curves versus time were taken to indicate a Mpemba effect. Such crossings, for initial temperatures above and below the glass transition and with subsequent relaxation below the glass transition temperature, can be seen on Figs. 34 and 35. An inverse Mpemba effect was also observed; in this case, all three temperatures were below the glass temperature of the model.
Figure 34: Mpemba effect in the spin-glass system starting from a paramagnetic phase, which is quenched below the critical temperature, $T_{c}=1.102$, to the spin-glass phase, with $T_{\rm final}=0.7\approx 0.64\ T_{c}$. The hot start, ($T_{1}=1.3$, $\xi_{1}=12$, $E_{1}(t=0)=-1.6428$), overtakes the two cold starts ($T_{2}=1.2$, $\xi_{2}=5$, $E_{2}(t=0)=-1.6714$) and ($T_{3}=1.2$, $\xi_{3}=8$, $E_{3}(t=0)=-1.6738$). The correlation length grows for temperatures close to $T_{c}$; thus, the hot start, since it has the longest correlation length, is the “closest” to the state the system is relaxing to. The order of the two crossings implies the importance of the correlation length. The error bars are depicted in the thickness of the lines. Source: Reprinted with permission from [59].
@@
Figure 35: (a) Mpemba effect in the spin-glass system below the glass transition, $T_{c}=1.102$, for several correlation lengths. The energy relaxation curves show that the coherence length plays a major role in the relaxation. (b) Relationship between energy and coherence length $\xi$. A thermal protocol is responsible for the vertical moves in (b) that produce the Mpemba effect. Source: Reprinted with permission from [59].